# Conditional Latent Space Molecular Scaffold Optimization for Accelerated Molecular Design

### v1

@@ 3.3 Scaffolds
Scaffolds Bemis & Murcko (1996) are the stable core structures within molecules that serve as the framework for chemical modifications in drug design. Scaffolds retain the essential biological activity of the molecule. They play a crucial role in molecular design by providing a foundation for chemical modifications aimed at optimizing properties like QED. Researchers often use scaffolds to systematically explore chemical variations Schreiber (2000), Welsch et al. (2010), which can lead to the discovery of new compounds with improved properties.
@@ 3.4 Problem Definition
We denote the scaffold, which is the base of the modification, as $\bm{S}$, and the modified molecule as $\bm{S}^{\prime}$. Our goal is to efficiently find the modification that maximizes the molecular property P which is evaluated by $f^{\text{BB}}(\bm{S}^{\prime})$, while keeping the difference between $\bm{S}$ and $\bm{S}^{\prime}$ small. Directly optimizing over all possible $\bm{S}^{\prime}$ to find the best modification that maximizes $f^{\text{BB}}(\bm{S\textquoteright})$ is impractical, as it involves high complexity and costly evaluations of $f^{\text{BB}}$.
@@
$\bm{z^{*}},p_{i}^{*}=\arg\max_{\bm{z}\in\mathcal{Z},p_{i}\in B(\mathbf{S})}f^{\text{BB}}(\bm{S^{{}^{\prime}}})=\arg\max_{\bm{z}\in\mathcal{Z},p_{i}\in B(\mathbf{S})}f^{\text{BB}}g(\bm{S}\oplus f^{\text{dec}}(\bm{z}),p_{i}),$ (2)
@@ 3.4.1 Controlling Molecular Similarity
Current modification-based methods often fail to account for how changes impact molecular similarity between the original scaffold $\bm{S}$ and the updated scaffold $\bm{S^{\prime}}$, or any structure in general. Adding substructures typically increases molecular weight, which can hinder real-world applicability, especially when exceeding 500 Daltons, as indicated by Lipinski’s Rule of Five Lipinski et al. (2001). Higher molecular weight compounds are more difficult to synthesize, making them less suitable for molecular design. Additionally, it is sometimes necessary to ensure that modifications result in only minor adjustments to avoid drastic changes.
@@ 4.1 Data Preparation
Our proposed method requires a uniquely tailored dataset because no existing dataset in the literature fully meets the specific needs of our approach. To create this dataset, we developed a BRICS (Breaking Retrosynthetically Interesting Chemical Substructures) Degen et al. (2008) based approach. BRICS is an algorithm designed to decompose organic molecules into smaller, synthetically feasible substructures by identifying breaking points within a molecule’s structure based on chemical retrosynthetic rules. These breaking points, also known as division points, are the connections between subgraphs within the molecule that BRICS identifies. The algorithm systematically breaks down a molecule $\bm{M}$ into $k$ substructures, Fig. 3 demonstrates this procedure.
The division points found by BRICS are of particular importance because they serve as both the points where the molecule is split into substructures and the potential bonding sites where these substructures can be reattached. Therefore, they provide us a valuable information about these substructures in terms of what kind of bonds they can form. This dual role makes the division points a crucial piece of information for our generative model. By preserving the properties and features of atoms at these division points, we capture the chemical context around the atom that dictates how substructures can bond with other parts of a molecule. The atomic features we consider are atom type, hybridization, valence, formal charge, degree, and ring membership. Detailed explanations of these features are provided in the Appendix A.1. These features, when used as condition vectors in our CVAE, are essential for guiding the generation of substructures that can successfully bond with the scaffold, as they provide a detailed understanding of the chemical environment at each division point, such as their atom types and if the selected atom has the capacity for forming additional bonds. Therefore, our dataset for CVAE training includes the substructures obtained via the BRICS algorithm and six different atomic features of the atom at their breaking points. The illustration of the BRICS algorithm, atomic feature extraction points, and their integration with CVAE is provided in Fig. 3.
@@ 4.2 Substructure CVAE
$\mathcal{L}(\bm{\phi},\bm{\theta};\mathbf{s},\bm{c})=\|\bm{s}-\mathbb{E}_{q_{\bm{\phi}}(\bm{z}|\bm{s},\bm{c})}[p_{\bm{\theta}}(\bm{s}|\bm{z},\bm{c})]\|^{2}+\beta\cdot D_{\text{KL}}(q_{\bm{\phi}}(\bm{z}|\bm{s},\bm{c})\|p(\bm{z})),$ (3)
where $q_{\bm{\phi}}(\bm{z}|\bm{s},\bm{c})$ is the encoder, $p_{\bm{\theta}}(\bm{s}|\bm{z},\bm{c})$ is the decoder, and $\beta$ balances the reconstruction and KL divergence terms Higgins et al. (2016). The condition vectors allow the CVAE to learn how different substructures interact with specific atomic environments, building a deeper understanding of their bonding behavior. Figure 3B demonstrates the input substructures that CVAE uses in its training, and Fig. 3C visualizes the CVAE architecture along with the condition vectors.
@@ 4.3 CLaSMO
In CLaSMO, we perform a targeted search in the latent space $\mathcal{Z}\in\mathbb{R}^{d}$ of the CVAE, where decodings from each point $\bm{z}$ represent a potential substructure to be added to the scaffold. The challenge is to modify the scaffold $\bm{S}$ by selecting a substructure and integrating it at an appropriate bonding point $p_{i}$, optimizing the desired molecular property $\mathcal{P}$. Therefore, the search space $\Omega$ we consider in our optimization tasks is defined as the Cartesian product of the latent space and the set of possible bonding points on the scaffold:
@@
However, although our main goal is optimizing the target property, we also aim to keep the similarity between $\bm{S}$ and $\bm{S^{\prime}}$ at a certain level. Therefore, in CLaSMO, we apply a similarity constraint using the Dice Similarity Dice (1945) metric to compare the input scaffold $\bm{S}$ with the modified molecule $\bm{S^{\prime}}$ before evaluating the BB function. In order to measure the similarity between $\bm{S}$ and $\bm{S^{\prime}}$, we use Morgan Fingerprints (Morgan, 1965, Rogers & Hahn, 2010). The Morgan fingerprint of a molecule is represented as a binary vector, where each bit indicates the presence or absence of a specific substructure within the molecule, making them effective and popular for comparing molecular similarities. Using these, the similarity between $\bm{S}$ and $\bm{S^{\prime}}$ is measured by computing the Dice Similarity of their Morgan fingerprint vectors111 Bajusz et al. (2015) discusses that the dice similarity is one of the best metrics to assess similarity between molecules using Morgan fingerprints., defined as:
@@
[…]
@@ 4.3.1 Kernel Design of CLaSMO
$k_{\text{cont}}(\bm{z},\bm{z}^{\prime})=\exp\left(-\frac{1}{2\ell^{2}}\|\bm{z}-\bm{z}^{\prime}\|^{2}\right),\quad k_{\text{cat}}(p_{i},p_{j})=\exp\left(-\frac{\delta_{p_{i},p_{j}}}{\ell}\right),$
where $k_{\text{cont}}(\bm{z},\bm{z}^{\prime})$ is an RBF kernel, and $k_{\text{cat}}(p_{i},p_{j})$ measures the similarity between atoms in the molecule that are ready for additional bond via Kronecker delta function, measuring equality of atom positions, and $\ell$ is the lengthscale parameter. The combined kernel used in CLaSMO is then expressed as:
@@
7:         Identify available atoms in the scaffolds for bonding, $p_{j}\in B(\bm{S})$
@@ 5 Experiments
In this section, we first provide details about the training of our CVAE model. Next, we discuss the results from two different experimental settings, which are QED property optimization and docking score property optimization.
@@ 5.1 CVAE Training
For our data preparation, we used the QM9 dataset Ruddigkeit et al. (2012), Ramakrishnan et al. (2014), which contains 130,000 small molecules. QM9 was chosen for its simplicity and suitability for our task, as it consists of smaller molecules compared to other well-known datasets like ZINC Irwin & Shoichet (2005), Irwin et al. (2020). Using our data preparation strategy, we extracted 18,706 unique substructure pairs along with their atomic environment features to train our CVAE. Most of the substructures obtained from QM9 through our method are under 100 Daltons, making them ideal for training a model focused on generating small substructures to modify the input scaffold. Since our goal is to optimize the scaffold while maintaining a high degree of similarity in the modifications, using such a dataset allows CVAE to learn to generate small substructures, which is crucial for achieving our goals. Of the 18,706 instances, 80% were used for model training, with the remaining data allocated for testing and validation. We represented the molecules using SELFIES Krenn et al. (2020), a string-based molecular representation, in the form of one-hot encoding matrices. Our CVAE architecture consists of three fully connected layers in the encoder and three GRU layers in the decoder.
@@
Given the simplicity of the dataset, we determined that a 2-dimensional latent space was sufficient to achieve over 99% reconstruction accuracy on the test set. This lower-dimensional space also enhances the performance of BO, which typically excels in smaller-dimensional spaces. Additionally, to maintain this simplicity, as explained in Section 4.2.1, we generated a 2-dimensional embedding for the condition vectors using an Autoencoder model, which is trained with fully connected layers in both the encoder and decoder, achieved 93% reconstruction accuracy for the condition vectors. Prior to CVAE training, we obtained the condition vector embeddings via the encoder of this VAE model, and used them during CVAE training. As a result, the CVAE was trained with both a 2-dimensional latent space for the substructures and 2-dimensional condition vectors, effectively balancing simplicity and optimization performance.
5.2 QED Optimization
In this section, we present the results of our QED optimization experiments, where we used RDKit Landrum (2010) to calculate QED values. The QED metric, defined between 0 and 1, inherently limits the range of potential improvements, making optimization within this closed range particularly challenging. To further complicate this task, we imposed similarity constraints to ensure the optimized molecules remain close to their original scaffolds, by running the CLaSMO algorithm with $\tau$ values $\tau\in[0,0.25,0.5,0.6]$. For each threshold, CLaSMO was run for 100 iterations on 100 input scaffolds, where their whole molecules are sampled randomly from the ZINC250K dataset Gómez-Bombarelli et al. (2018). The search domain for each latent dimension in $\mathcal{Z}$ is set to $[-6,6]$. The GP model is trained using 100 training instances. Such an experimental setting is designed to demonstrate the optimization capabilities of CLaSMO at changing similarity thresholds, within the limited range of QED optimization tasks.
Figure 4: Distribution of QED values of input scaffolds and CLaSMO optimization results at various similarity thresholds, illustrating the shift towards higher QED values at each similarity level. (A) $\tau=0$, showing the greatest improvement due to no similarity constraints. (B) $\tau=0.25$, demonstrating high QED gains. (C) $\tau=0.5$ and (D) $\tau=0.6$ are still able to demonstrate significant improvements while enforcing a balance between optimization and similarity.
Initially, the average QED score of the input scaffold molecules was 0.5876, with a maximum QED of 0.8902. After optimization, CLaSMO with no similarity constraint ($\tau=0$) achieved the highest QED score of 0.9480 and an average QED of 0.6778, representing a mean improvement rate of 21.43% [^2] Mean improvement rates calculated as the average of the relative improvements obtained by CLaSMO for each input scaffolds. and a maximum improvement rate of 81.17%. As the similarity constraint was tightened, the performance remained competitive. With $\tau=0.25$, the maximum QED score was 0.9464, with an average QED of 0.6715, leading to a mean improvement rate of 19.14%. Further increasing the similarity threshold to $\tau=0.50$, the maximum QED score was 0.9437, and the average QED was 0.6602, yielding a mean improvement rate of 16.08%. For $\tau=0.60$, the maximum QED score was 0.9402, and the average QED was 0.6434, resulting in a mean improvement rate of 11.70%. The results clearly demonstrate that CLaSMO consistently enhances the QED values across a range of similarity thresholds, effectively balancing the trade-off between maximizing QED and preserving structural similarity. Notably, even under more stringent similarity constraints, CLaSMO achieves significant improvements over the initial QED values, highlighting its robust optimization capabilities. This consistent performance across different thresholds underscores CLaSMO’s versatility and effectiveness in optimizing QED, making it a powerful tool for scaffold-based molecular optimization tasks. Table 1 summarizes these findings. These results were achieved with only 100 iterations per input scaffold, demonstrating the sample efficiency of our approach. Additionally, Fig. 4 demonstrates two example sets of scaffold optimization results from the CLaSMO experiments, in which it can be observed that our algorithm finds a molecule with a QED score of 0.948 by very limited additions, keeping the similarity between the input and optimized molecule at the high level.
Table 1: Summary of QED optimization results: Max and mean QED values from input scaffolds and CLaSMO experiments with various $\tau$ levels, as well as mean and max improvement rates over QED values of input scaffold.
Metric Input $\bm{\tau=0}$ $\bm{\tau=0.25}$ $\bm{\tau=0.50}$ $\bm{\tau=0.60}$
Max QED 0.8902 0.9480 0.9464 0.9437 0.9402
Mean QED 0.5876 0.6778 0.6715 0.6602 0.6434
Mean Imp. - 21.43% 19.14% 16.08% 11.70%
Max Imp. - 81.17% 114.57% 74.74% 72.41%

### v2

@@ 3.3 Scaffolds
Scaffolds (Bemis & Murcko, 1996) are the stable core structures within molecules that serve as the framework for chemical modifications in drug design. Scaffolds retain the essential biological activity of the molecule. They play a crucial role in molecular design by providing a foundation for chemical modifications aimed at optimizing properties like QED. Researchers often use scaffolds to systematically explore chemical variations (Schreiber, 2000, Welsch et al., 2010), which can lead to the discovery of new compounds with improved properties.
@@ 3.4 Problem Definition
We denote the scaffold, which is the base of the modification, as $\bm{S}$, and the modified molecule as $\bm{S}^{\prime}$. Our goal is to efficiently find the modification that maximizes the molecular property P which is evaluated by $f^{\text{BB}}(\bm{S}^{\prime})$, while keeping the difference between $\bm{S}$ and $\bm{S}^{\prime}$ small. Directly iterating over all possible $\bm{S}^{\prime}$ to find the best modification that maximizes $f^{\text{BB}}(\bm{S’})$ is impractical, as it involves high complexity and costly evaluations of $f^{\text{BB}}$.
@@
$\bm{z^{*}},p_{i}^{*}=\arg\max_{\bm{z}\in\mathcal{Z},p_{i}\in B(\mathbf{S})}f^{\text{BB}}(\bm{S^{{}^{\prime}}})=\arg\max_{\bm{z}\in\mathcal{Z},p_{i}\in B(\mathbf{S})}f^{\text{BB}}(g(\bm{S}\oplus f^{\text{dec}}(\bm{z}),p_{i})),$ (2)
@@ 3.4.1 Controlling Molecular Similarity
Current modification-based methods often fail to account for how changes impact molecular similarity between the original scaffold $\bm{S}$ and the updated scaffold $\bm{S^{\prime}}$, or any structure in general. Adding substructures typically increases molecular weight, which can hinder real-world applicability, especially when exceeding 500 Daltons, as indicated by Lipinski’s Rule of Five (Lipinski et al., 2001). Higher molecular weight compounds are more difficult to synthesize, making them less suitable for molecular design. Additionally, it is sometimes necessary to ensure that modifications result in only minor adjustments to avoid drastic changes.
@@ 4.1 Data Preparation
Our proposed method requires a uniquely tailored dataset because no existing dataset in the literature fully meets the specific needs of our approach. To create this dataset, we developed a BRICS (Breaking Retrosynthetically Interesting Chemical Substructures) (Degen et al., 2008) based approach. BRICS is an algorithm designed to decompose organic molecules into smaller, synthetically feasible substructures by identifying breaking points within a molecule’s structure based on chemical retrosynthetic rules. These breaking points, also known as division points, are the connections between subgraphs within the molecule that BRICS identifies. The algorithm systematically breaks down a molecule $\bm{M}$ into $k$ substructures, Fig. 3 demonstrates this procedure. BRICS is particularly well-suited for this task because it ensures that the generated substructures are synthetically feasible and chemically valid, making it ideal for scaffold-based molecular optimization. Unlike other decomposition methods, BRICS explicitly adheres to retrosynthetic rules, ensuring compatibility with real-world chemical synthesis.
The division points found by BRICS are of particular importance because they serve as both the points where the molecule is split into substructures and the potential bonding sites where these substructures can be reattached. Therefore, they provide us a valuable information about these substructures in terms of what kind of bonds they can form. This dual role makes the division points a crucial piece of information for our generative model. By preserving the properties and features of atoms at these division points, we capture the chemical context around the atom that dictates how substructures can bond with other parts of a molecule. The atomic features we consider are atom type, hybridization, valence, formal charge, degree, and ring membership. Detailed explanations of these features are provided in the Appendix A.1. These features, when used as condition vectors in our CVAE, are essential for guiding the generation of substructures that can successfully bond with the scaffold, as they provide a detailed understanding of the chemical environment at each division point, such as their atom types and if the selected atom has the capacity for forming additional bonds. Therefore, our dataset for CVAE training includes the substructures obtained via the BRICS algorithm and six different atomic features of the atom at their breaking points. These substructures are represented using a string-based representation method referred to as SELFIES Krenn et al. (2020). The illustration of the BRICS algorithm, atomic feature extraction points, and their integration with CVAE is provided in Fig. 3.
@@ 4.2 Substructure CVAE
$\textbf{J}(\bm{\phi},\bm{\theta};\mathbf{s},\bm{c})=\|\bm{s}-\mathbb{E}_{q_{\bm{\phi}}(\bm{z}|\bm{s},\bm{c})}[p_{\bm{\theta}}(\bm{s}|\bm{z},\bm{c})]\|^{2}+\beta\cdot D_{\text{KL}}(q_{\bm{\phi}}(\bm{z}|\bm{s},\bm{c})\|p(\bm{z})),$ (3)
where $q_{\bm{\phi}}(\bm{z}|\bm{s},\bm{c})$ is the encoder, $p_{\bm{\theta}}(\bm{s}|\bm{z},\bm{c})$ is the decoder, and $\beta$ balances the reconstruction and KL divergence terms (Higgins et al., 2016). The condition vectors allow the CVAE to learn how different substructures interact with specific atomic environments, building a deeper understanding of their bonding behavior. Figure 3B demonstrates the input substructures that CVAE uses in its training, and Fig. 3C visualizes the CVAE architecture along with the condition vectors.
@@ 4.3 CLaSMO
In CLaSMO, upon training of the CVAE, we perform a targeted search in the latent space $\mathcal{Z}\in\mathbb{R}^{d}$ of the CVAE, where decodings from each point $\bm{z}$ represent a potential substructure to be added to the scaffold. The challenge is to modify the scaffold $\bm{S}$ by selecting a substructure and integrating it at an appropriate bonding point $p_{i}$, optimizing the desired molecular property $\mathcal{P}$. Therefore, the search space $\Omega$ we consider in our optimization tasks is defined as the Cartesian product of the latent space and the set of possible bonding points on the scaffold:
@@
However, although our main goal is optimizing the target property, we also aim to keep the similarity between $\bm{S}$ and $\bm{S^{\prime}}$ at a certain level. Therefore, in CLaSMO, we apply a similarity constraint using the Dice Similarity (Dice, 1945) metric to compare the input scaffold $\bm{S}$ with the modified molecule $\bm{S^{\prime}}$ before evaluating the BB function. In order to measure the similarity between $\bm{S}$ and $\bm{S^{\prime}}$, we use Morgan Fingerprints (Morgan, 1965, Rogers & Hahn, 2010). The Morgan fingerprint of a molecule is represented as a binary vector, where each bit indicates the presence or absence of a specific substructure within the molecule, making them effective and popular for comparing molecular similarities. Using these, the similarity between $\bm{S}$ and $\bm{S^{\prime}}$ is measured by computing the Dice Similarity of their Morgan fingerprint vectors [^2] Bajusz et al. (2015) discusses that the dice similarity is one of the best metrics to assess similarity between molecules using Morgan fingerprints., defined as:
@@
$\text{subject to:}\quad\text{DICE}(\bm{S},\bm{S^{\prime}})\geq\tau,$
@@ 4.3.1 Kernel Design of CLaSMO
$k_{\text{cont}}(\bm{z},\bm{z}^{\prime})=\exp\left(-\frac{1}{2\ell^{2}_{1}}\|\bm{z}-\bm{z}^{\prime}\|^{2}\right),\quad k_{\text{cat}}(p_{i},p_{j})=\exp\left(-\frac{\delta_{p_{i},p_{j}}}{\ell_{2}}\right),$
where $k_{\text{cont}}(\bm{z},\bm{z}^{\prime})$ is an RBF kernel, and $k_{\text{cat}}(p_{i},p_{j})$ measures the similarity between atoms in the molecule that are ready for additional bond via Kronecker delta function, measuring equality of atom positions, $\ell_{1}$ and $\ell_{2}$ are the lengthscale parameters for the continuous and categorical components, respectively. The combined kernel used in CLaSMO is then expressed as:
@@
7:   Identify available atoms in the scaffolds for bonding, $p_{j}\in B(\bm{S})$ [^2] This identification is performed by finding atoms within the molecule that have additional capacity to form bonds. For example, a carbon atom can form up to four bonds, and if it has only three bonds within the molecule, it will be identified as available for bonding.
@@ 5 Experiments
In this section, we first describe the training procedure for our CVAE model (Section 5.1). We start from our dataset selection procedure, and introduce the details of our model architecture. We then evaluate the performance of CLaSMO across 20 molecular property optimization tasks Gao et al. (2022), which span a range of objectives including rediscovery of known compounds, similarity-based optimization, multi-property optimization, and improvements in drug-likeness.
To assess robustness across different optimization regimes, each task is repeated under three levels of molecular similarity constraints between the input and optimized molecules, defined by Dice similarity thresholds of 0.50, 0.25, and 0. This results in a total of 60 optimization settings. Higher thresholds enforce stricter structural similarity, enabling controlled, local edits, while lower thresholds allow for broader exploration of chemical space. We also evaluate the synthetic and retrosynthetic accessibilities of the generated molecules to assess their feasibility from a chemical synthesis and retrosynthetic planning perspective.
Section 5.2 presents the results of these optimization experiments and the corresponding synthetic and retrosynthetic accessibility analysis. In Section 5.3, we provide an ablation study to examine the impact of conditional generation. Finally, Section 5.4 discusses the extension of CLaSMO to a human-in-the-loop optimization setting.
@@ 5.1 CVAE Training
In this section, we start from the dataset selection procedure and then provide model training details.
5.1.1 Dataset Selection
Our motivation for using a BRICS-curated dataset is twofold. First, BRICS decomposition enables interpretable fragmentation of molecules by producing chemically meaningful substructures with explicit connection points. These points can be directly associated with atomic features such as atom type, degree, and hybridization, which allows us to condition the generative model on the chemical context at each attachment site. Second, since our overall objective is to develop a generative model that performs small, targeted modifications to existing molecules, it is essential to train on substructures that reflect realistic, low-molecular-weight chemical fragments. This encourages the model to propose minimal, context-aware edits to molecular scaffolds during inference.
To this end, we evaluated two widely used open-source molecular datasets: QM9 (Ruddigkeit et al., 2012, Ramakrishnan et al., 2014) and ZINC250K (Gómez-Bombarelli et al., 2018). We applied BRICS decomposition to both datasets and analyzed the resulting fragment distributions with a focus on their suitability in our setting. As shown in Figure 4, BRICS fragments derived from QM9 are compact and chemically consistent, with all fragments under 105 Da and most well below 100 Da. In contrast, fragments from ZINC250K exhibit greater chemical complexity and structural variability, with a substantial portion exceeding 100 Da in molecular weight. Representative examples of fragments from both datasets are also presented in Figure 4.
Based on this analysis, we selected QM9 as our primary dataset for training the CVAE model. QM9 contains approximately 130,000 small organic molecules and, when processed via BRICS, yields 18,706 unique substructure pairs along with their atomic environment features. The relatively small size and chemical simplicity of these fragments align well with our modeling objective of generating small, synthetically accessible modifications to existing scaffolds. Additionally, the compact nature of QM9 fragments eliminates the need for further filtering, making the dataset highly consistent and easy to work with. This makes QM9 particularly suitable for our use case, where the goal is to preserve high similarity to the input scaffold while introducing meaningful yet minimal structural improvements [^4] In Appendix A.2, we explore two possible extensions: one where the CVAE is trained exclusively on BRICS fragments derived from the ZINC250K dataset, and another where it is trained on a mixture of QM9 fragments and size-compatible fragments filtered from ZINC250K.
5.1.2 Model Training
Among the 18,706 instances, 80% were used for model training, with the remaining data allocated for testing and validation. We represented the molecules using SELFIES (Krenn et al., 2020), a string-based molecular representation, in the form of one-hot encoding matrices. Our CVAE architecture consists of three fully connected layers in the encoder and three GRU layers in the decoder. Given the simplicity of the dataset, we determined that a 2-dimensional latent space was sufficient to achieve over 90% reconstruction accuracy on the test set, where training is conducted using an early stopping condition along with a learning rate scheduler. This lower-dimensional space also enhances the performance of BO, which typically excels in smaller-dimensional spaces. See Appendix A.3 for further details on the selection of hyperparameters in CLaSMO and in training the CVAE model.
@@ 5.1.2 Model Training
Additionally, to maintain this simplicity, as explained in Section 4.2.1, we generated a 2-dimensional embedding for the condition vectors using an Autoencoder model, which is trained with fully connected layers in both the encoder and decoder, achieved 93% reconstruction accuracy for the condition vectors. Prior to CVAE training, we obtained the condition vector embeddings via the Autoencoder model, and used them during CVAE training. As a result, the CVAE was trained with both a 2-dimensional latent space for the substructures and 2-dimensional condition vectors, effectively balancing simplicity and optimization performance. Further discussion on this design choice is provided in Appendix A.4.
Figure 4: Distribution of BRICS fragment molecular weights from the QM9 and ZINC250K datasets, along with representative examples from each. BRICS fragments from the QM9 dataset generally have lower molecular weights, reflecting its focus on smaller molecules, while the ZINC250K dataset contributes a broader range of fragment sizes.
5.2 Molecular Optimization Experiments
To comprehensively evaluate the sample-efficiency of our methodology, we applied CLaSMO to the sample-efficiency benchmark proposed by Gao et al. (2022), where we use 20 diverse molecular optimization tasks. These tasks span various objectives, including a broad range of pharmaceutically-relevant oracle functions, and challenging multi-property optimization (MPO) scenarios (Brown et al., 2019). Instead of limiting the evaluation to narrowly defined tasks, this broad selection enables a more comprehensive assessment of each method’s generalizability and adaptability across chemically and functionally diverse scenarios.
Here, sample-efficiency refers to the efficiency with respect to the number of evaluations of the black-box function for the target optimization or editing task. For instance, when evaluating molecules through wet-lab experiments, it is often the case that no more than about 100 evaluations can be conducted. Our proposed method, CLaSMO, is specifically designed to achieve such sample-efficient molecular editing. On the other hand, in the numerical experiments presented in this paper, we selected problems where the evaluation of the target black-box functions can be performed at relatively low computational cost, allowing for repeated and comprehensive comparisons between the proposed method and existing methods. Hereafter, we refer to these functions as oracle functions. Experiments were conducted with a fixed budget of 100 oracle evaluations per seed across 10 seeds, ensuring consistent comparisons. The choice of 100 oracle iterations (in contrast to the 10,000 iterations used in Gao et al. (2022)) reflects our focus on developing a methodology tailored for resource-constrained scenarios, which are more representative of real-world applications.
CLaSMO was benchmarked against six widely used baseline methods—Smiles-GA (Brown et al., 2019), Graph-GA (Jensen, 2019), Stoned (Nigam et al., 2021), MolDQN (Zhou et al., 2019a), MARS (Xie et al., 2021), and SynFlowNet (Cretu et al., 2024)—using the same set of 100 scaffolds for all methods as starting molecules. To assess optimization performance across scaffolds of varying sizes, we randomly sampled input scaffolds from both the QM9 and ZINC250K datasets. Appendix A.5 provides detailed statistics on their molecular weight and property value distributions, along with visualizations of all 100 scaffolds used in the evaluation. As explained below, different similarity constraints were applied, and open-source implementations of all baseline methods were used to evaluate their ability to perform molecular optimization while preserving the given scaffold. We note that some of these baseline methods were not designed for sample-efficient settings, and therefore may not perform well under the problem settings aimed at sample-efficient molecular editing, which is the focus of this study. [^5] Among the competing approaches, Graph-GA, Smiles-GA, and Stoned lack a mechanism to ensure preservation of the given scaffold. We implemented a hyperparameter optimization for these genetic algorithm-based approaches to find the best balance for exploration of the search space and scaffold-preserving capabilities. Please refer to Appendix A.6 for further details on experimental setting.
5.2.1 Oracle Definitions and Grouping
To facilitate clearer comparisons and highlight performance patterns, we group the 20 oracles used in our experimental setting into four task-oriented categories. The groups and their associated oracle functions are as follows:
  • Group 1: Similarity and Rediscovery Tasks This group comprises oracles that either assess molecular similarity to a known drug or target the rediscovery of a known molecule. The oracles in this group include: albuterol_similarity, celecoxib_rediscovery, mestranol_similarity, thiothixene_rediscovery, troglitazone_rediscovery.
  • Group 2: Multi-Property Optimization (MPO) Tasks These involve optimizing multiple molecular properties simultaneously. The oracles in this group include: osimertinib_mpo, perindopril_mpo, ranolazine_mpo, sitagliptin_mpo, zaleplon_mpo.
  • Group 3: Structural Modification Tasks This group encompasses tasks that evaluate a model’s ability to generate structurally diverse molecules while preserving specific chemical features. The hopping tasks challenge models to modify the molecule while aiming to maintain or enhance biological activity despite structural changes. In contrast, the isomers tasks focus on generating different structural arrangements (isomers) of a molecule that share the same molecular formula but differ in connectivity or spatial orientation. Unlike similarity-based tasks that prioritize closeness to a reference molecule, these tasks emphasize structural innovation and diversity, assessing a model’s capacity to explore novel chemical spaces while retaining desired properties. The oracles in this group include: deco_hop, scaffold_hop, isomers_c7h8n2o2, isomers_c9h10n2o2pf2cl.
  • Group 4: Activity and Benchmark Tasks This group includes oracles that predict biological activity or are part of broader benchmarking suites for drug-likeness and diversity. The oracles in this group include: drd2, gsk3b, median1, median2, qed, valsartan_smarts.
Further details and descriptions of these oracles are provided in Appendix A.7.666See also the Therapeutics Data Commons for more detail: https://tdcommons.ai/functions/oracles/.
5.2.2 Property Optimization Results
In this section, we discuss the performance of CLaSMO and benchmark methodologies under three similarity-constraint regimes. In the high constraint setting ($\tau=0.50$), the optimized molecule must have a Dice similarity of at least 0.50 to the input scaffold. In the moderate constraint setting ($\tau=0.25$), a similarity of at least 0.25 is required, allowing for more flexibility in modifying the scaffold. In these settings, even if the target property can be substantially improved, the result will not be accepted if the similarity falls below the specified threshold. Lastly, in the no constraint setting ($\tau=0$), there are no restrictions on similarity, and any substructure additions to the scaffold are allowed [^7] See Appendix A.8 for a further discussion of the rationale behind the chosen similarity constraint thresholds..
Optimization with High Constraints (Similarity Threshold $\tau$ = 0.50)
Figure 5 shows the performance of the methods when the similarity constraint is very strict.
  • In the Similarity and Rediscovery Tasks, CLaSMO is consistently the top performer, providing the top performance from 5 tasks out of 5.
  • Within the MPO Tasks, CLaSMO and MARS methodologies are the two top performers. MARS, a method that is specifically designed for MPO settings, outperforms CLaSMO in two out of 5 MPO objectives, and provides a similar performance in osimertinib_mpo task.
  • For the Structural Tasks, CLaSMO gives the best results in deco_hop, scaffold_hop and isomers_c7h8n2o2 while MARS achieves a similar result with CLaSMO in the remaining isomers_c9h10n2o2pf2cl task.
  • In the Activity and Benchmark Tasks, CLaSMO records the highest values for drd2, median1 and qed, while losing to MARS in gsk3b and median2. None of the methodologies provided an improvement in valsartan_smarts task.
Overall, under high constraints, CLaSMO proves to be the best approach, providing the highest Top-10 average scores in 13 out of 20 property optimization tasks.
Figure 5: Comparison of Top-10 average values obtained at each iteration for CLaSMO, MARS, Smiles-GA, SynFlowNet, Graph-GA, MolDQN, and Stoned using constraint optimization setting, where Dice Similarity threshold $\tau=0.50$. CLaSMO identifies molecules with a higher Top-10 average score after 100 iterations in most of the optimization tasks. Among benchmark methodologies, MARS is competitive, followed by Graph-GA and Stoned.
Optimization with Moderate Constraints (Similarity Threshold $\tau$ = 0.25)
Figure 6 summarizes the results under a moderate similarity constraint, which allows for more chemical modification while still preserving significant scaffold features.
  • In the Similarity and Rediscovery Tasks, CLaSMO maintains its lead, where it achieves the highest Top-10 average scores in 4 out of 5 tasks, losing to MARS only in thiothixene_rediscovery.
  • Among the MPO Tasks, MARS is the top performer, where it provides the highest Top-10 average scores in 3 out of 5 tasks. CLaSMO outperforms MARS in only zaleplon_mpo task, and performs similarly with MARS in osimertinib_mpo task.
  • In the Structural Tasks, CLaSMO gives the best results in all of the 4 tasks in this group.
  • In the Activity and Benchmark Tasks, CLaSMO records the highest values for median1, median2 and qed, while losing to MARS in drd2 and gsk3b. None of the methodologies provided an improvement in valsartan_smarts task.
Thus, even with moderate flexibility, CLaSMO is able to harness the allowed modifications to deliver the best overall results, providing the best results in 13 optimization tasks out of 20.
Figure 6: Comparison of Top-10 average values obtained at each iteration for CLaSMO, MARS, Smiles-GA, SynFlowNet, Graph-GA, MolDQN, and Stoned using constraint optimization setting, where Dice Similarity threshold $\tau=0.25$. CLaSMO identifies molecules with a higher Top-10 average score after 100 iterations in most of the optimization tasks. Among benchmark methodologies, MARS is competitive, followed by Graph-GA and Stoned.
Optimization with No Constraints (Similarity Threshold $\tau$ = 0)
When no similarity constraints are imposed (see Fig. 7), methods only need to keep the scaffold, and all additions to the scaffold are allowed.
  • Similarity and Rediscovery Tasks, CLaSMO maintains its lead, where it achieves the highest Top-10 average scores in 4 out of 5 tasks, losing to MARS only in thiothixene_rediscovery.
  • In the MPO Tasks, MARS gets the best results in 3 out of 5 objectives, and CLaSMO finds the best results in zaleplon_mpo task and performs similary as MARS in osimertinin_mpo task.
  • In the Structural Tasks, CLaSMO gives the best results in all of the 4 tasks in this group.
  • In the Activity and Benchmark Tasks, CLaSMO records the highest values for median1 and qed, while losing to MARS in drd2, gsk3b and median2. None of the methodologies provided an improvement in valsartan_smarts task.
Figure 7: Comparison of Top-10 average values obtained at each iteration for CLaSMO, MARS, Smiles-GA, SynFlowNet, Graph-GA, MolDQN, and Stoned, without any constraint threshold. CLaSMO identifies molecules with a higher Top-10 average score after 100 iterations in most of the optimization tasks. Among benchmark methodologies, MARS is highly competitive, followed by Graph-GA and Stoned.